Parallel plates with an area of 72 cm^2 are separated by 2.6 cm of air and are storing 36 uC of charge. What is the potential difference across the plates? (If you happen to know equations involving capacitance, don't use those yet.) (The final answer is 1.47x10^7 V)

3. Parallel plates with an area of 72 cm2 are separated by 2.6 cm of air and are storing 36 uC of charge. What is the potential difference across the plates? (If you happen to know equations involving capacitance, dont use those yet) [1.47x10 V]
Answer #1

capacitance = Aeo / d

where A = Area , d = separation

C = 72*10^-4*8.85*10^-12 / 0.026

=2.45*10^-12 F

= 2.45 pF

V = Q/C

= 36*10^-6 / (2.45*10^-12) = 1.47*10^7 V